Exodus Limousine Company has $1,000 par value bonds outstanding at 10 percent interest. The bonds will mature in 50 years. Use Appendix B and Appendix D for an approximate answer but calculate your final answer using the formula and financial calculator methods. Compute the current price of the bonds if the percent yield to maturity is: (Do not round intermediate calculations. Round your final answers to 2 decimal places. Assume interest payments are annual.)

A. 5%

B. 15%
Answer #1

Interest rate 5%:

Coupon = 0.1 * 1000 = 100

Current price = Coupon * [1 - 1 /(1 +r)n] / r + FV / (1 + r)n

Current price = 100 * [1 - 1 /(1 +0.05)50] / 0.05 + 1000 / (1 + 0.05)50

Current price = 100 * [1 - 0.087204] /0.05 + 87.203727

Current price = 100 * 18.25592 + 87.203727

Current price = $1,912.80

Interest rate 15%:

Coupon = 0.1 * 1000 = 100

Current price = Coupon * [1 - 1 /(1 +r)n] / r + FV / (1 + r)n

Current price = 100 * [1 - 1 /(1 +0.15)50] / 0.15 + 1000 / (1 + 0.15)50

Current price = 100 * [1 - 0.000923] /0.15 + 0.922801

Current price = 100 * 6.660515 + 0.922801

Current price = $666.97